# Generating variable $\hbar$ and $c$ via Fujii–Wetterich model in curved spacetimes

### v1

@@ V. Resolving the age problem, bypassing DE—The flat $\Lambda$CDM model gives the age formula
VII. Physical intuition of late-time acceleration via VSL—Consider two supernovae, $A$ and $B$, located at distances $d_{A}$ and $d_{B}$ from the Earth, with $d_{B}=2\,d_{A}$. In standard cosmology, their redshift values $z_{A}$ and $z_{B}$ are related as $z_{A}\approx 2\,z_{A}$ (to first-order approximation). However, this relationship breaks down in our VSL cosmology. Since $c\propto a^{-1/2}$, light traveled faster in the distant past (when the cosmic factor $a\ll 1$) than in the more recent epoch (when $a\lesssim 1$). Therefore, photons emitted from supernova $B$ could cover twice the distance in less than double the time required by photons emitted from supernova $A$. Having spent less time in transit than predicted by standard cosmology, the $B$–photons experience less cosmic expansion than expected, resulting in a lower redshift such that:
@@
This occurs because the $C$–photons traveled faster at the beginning of their journey toward Earth and must originate from a farther distance (thus appearing fainter than expected) to accumulate sufficient redshift. In either case, the VSL-based Hubble diagram exhibits a steeper upward slope in the high-$z$ section.
An alternative explanation for late-time acceleration is a decrease in the speed of light in the intergalactic void as the universe expands. In summary, VSL cosmology can account—qualitatively and quantitatively—for the distance modulus excess observed in high-$z$ SnIa without the need for dark energy. This approach overcomes the fine-tuning and coincidence problems associated with the concordance $\Lambda$CDM model. Furthermore, it bypasses the enigmatic nature of dark energy and aligns with BDRS’s 2003 groundbreaking analysis of the CMB power spectrum which also negates the dark energy hypothesis [54].
VIII. Resolving the $H_{0}$ tension via an astronomical origin—In recent years, interest in the $H_{0}$ tension has intensified [60]. One popular approach to address this issue involves treating $H_{0}$ as a “running” value, meaning that its value depends on the redshift of the data used to deduce it; e.g., see Ref. [61].
Our VSL cosmology naturally produces a “running” $H_{0}$ through the function $F(z)$. As mentioned in Point I, $F(z)$ represents the change in the typical size of galaxies as they evolve in an expanding intergalactic void (see also Ref. [14] for clarifications). From the modified formula (58), we can define the “running” value for $H_{0}$ as
@@
Using (54) and the value $F_{\infty}=0.93$ obtained from the fit in Point II, the computed “running” $H_{0}(z)$ is depicted as dashed curves in Fig. 3. At the time of recombination, the “running” value for $H_{0}$ would be $H_{0}\,F_{\infty}=47.2\times 0.93=43.9$, representing a $7\%$ reduction from its value estimated using low-redshift SnIa. The function $F(z)$ hence shed important light on the $H_{0}$ tension. Furthermore, in [55] Hunt and Sarkar obtained $H_{0}\approx 43.5$ from the CMB power spectrum using a supergravity-based multiple inflation scenario. Our result of $H_{0}(z\gg 1)\approx 43.9$ is thus in agreement with that of the Hunt–Sarkar team.
Importantly, the function $F(z)$ is a well-defined and physically motivated concept: it models the evolution of galaxies in an expanding intergalactic void. Consequently, the “running” $H_{0}$ in our approach is rooted in an astronomical origin, rather than a cosmic origin.
IX. An infinite cosmological horizon—In the generalized FLRW metric (50), the cosmological horizon at the recombination time $t_{rec}$ is given by
@@
III Conclusions
The action given in Eqs. (6), (7), and (12) describes a generally covariant theory of gravitation and matter, which also possesses dilatation symmetry. A notable feature is the non-minimal coupling of matter with gravity, expressed through the coupling of the Higgs field $\Phi$ with a background ‘scalon’ field $\chi$ in the form $\chi^{2}\,\Phi^{2}$. Crucially, the (dimensionful) quantum of action $\hbar$ and speed of light $c$ are absent in this adimensional action. Previously, Fujii and Wetterich utilized this action to generate variable particle masses and (Newton) gravitational ‘constant’ [8, 9, 10, 11, 12]. In Section I of this paper, we developed an alternative mechanism to construct (variable) $\hbar$ and $c$ from the background scalon field $\chi$. Our approach can be generalized when the matter Lagrangian (7) is extended to cover the Glashow–Weinberg–Salam (GWS) model of the $SU(3)\times SU(2)\times U(1)$ gauge group [13].
In our framework, each open set vicinity of any given point $x^{*}$ on the spacetime manifold is equipped with a replica of the GWS action for matter fields (fermions, gauge vector bosons, and Higgs scalar bosons), operating with its own effective values of $\hbar_{*}$ and $c_{*}$. Lorentz symmetry holds locally within each open set. These GWS replicas are isomorphic (i.e. structurally identical), differing only in the values of $\hbar_{*}$ and $c_{*}$, which are determined by the prevailing value of the background scalon field $\chi(x^{*})$ per $\hbar_{*}=\chi^{-1/2}(x^{*})$ and $c_{*}=\chi^{1/2}(x^{*})$. Crucially, within each open set, the length and time scales of physical processes are governed by the scalon field as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$, resulting in an anomalous anisotropic relation between the rate of clocks and the length of rods, $\tau^{-1}\propto l^{-3/2}$. This novel relation enables us to predict a new “time dilation effect of the Third kind”, with further details presented in Ref. [16].
Our construction hence fulfills Einstein’s original—albeit long overlooked—vision in incorporating the local validity of Lorentz symmetry, the variation of speed of light in curved spacetimes, and the geometric nature of gravity [18].
The anisotropic relation $\tau^{-1}\propto l^{-3/2}$ mentioned above gives rise to new physics and phenomenology. Section II of our paper presented key findings on the phenomenological consequences of our mechanism in the context of cosmology, particularly addressing late-time cosmic acceleration via the variation in $c$ while eliminating the need for dark energy. Remarkably, our results revive a crucial—but previously overlooked—discovery made by Blanchard, Douspis, Rowan-Robinson, and Sarkar in 2003 which circumvented the requirement for dark energy in their analysis of the CMB power spectrum [54].
@@ Acknowledgements.
I wish to thank Clifford Burgess for his instrumental feedback, which greatly sharpened the exposition of this paper. I also thank Anne-Christine Davis and Sergey Sibiryakov for their valuable insights. The constructive comments from Robert Mann, Tiberiu Harko, Tessa Baker, John Moffat, Niayesh Afshordi, Alexander Kamenshchik, Bertrand Chauvineau, Radosław Kycia, Péter Ván, Lorenzo Gavassino, and Gregory Horndeski are gratefully acknowledged.
@@
In our mechanism, the electron mass $m$ and charge $e$ are dimensionless quantities derived from the parameters $f$, $\mu$, $\lambda$, and $\alpha$ via Eq. (19). The scalon field $\chi$ is the single entity that carries a scale, which then serves as the only unit available. Consequently, all physical quantities can be expressed exclusively in terms of the magnitude of $\chi$. For instance, $\hbar$ and $c$ have the units of $\chi^{-1/2}$ and $\chi^{1/2}$, respectively, whereas $G$ is dimensionless. The rest mass energy $E=mc^{2}$ has the unit of $\chi$, owing to $m$ being dimensionless and $c\propto\chi^{1/2}$. Similarly, the Coulomb potential of a point charge $e$, given by $e^{2}/r$, also carries the unit of $\chi$, by virtue of $e$ being dimensionless and $r\propto\chi^{-1}$. In general, a theory with dilatation symmetry would require only one ‘unit’ (associated with the scalon field $\chi$), rendering the canonical three-unit system (mass $M$, length $L$, time $T$) unnecessary.
@@
Our new unit system based on the trio $\{e_{0},\,m_{0},\,\chi_{*}\}$ deviates from the Planck $cGh$ unit system, which relies on the trio of constants $\{c,\,G,\,\hbar\}$. Our new unit system is akin to the unit system introduced by Hartree in 1928 [62], based on the electron charge $e$, mass $m$, and the Bohr radius $a_{B}$ (the typical size of the hydrogen atom in its ground state). In the Hartree system:
$\hbar$ $=e\sqrt{m\,a_{B}};\ \ \ c=\frac{e}{\alpha\sqrt{ma_{B}}}\,.$ (75)

### v2

@@ VI.6 Obtaining late-time acceleration via VSL, bypassing DE
VI.7 Physical intuition of late-time acceleration via VSL
Consider two supernovae, $A$ and $B$, located at distances $d_{A}$ and $d_{B}$ from the Earth, with $d_{B}=2\,d_{A}$. In standard cosmology, their redshift values $z_{A}$ and $z_{B}$ are related as $z_{A}\approx 2\,z_{A}$ (to first-order approximation). However, this relationship breaks down in our VSL cosmology. Since $c\propto a^{-1/2}$, light traveled faster in the distant past (when the cosmic factor $a\ll 1$) than in the more recent epoch (when $a\lesssim 1$). Therefore, photons emitted from supernova $B$ could cover twice the distance in less than double the time required by photons emitted from supernova $A$. Having spent less time in transit than predicted by standard cosmology, the $B$–photons experience less cosmic expansion than expected, resulting in a lower redshift such that:
@@ VI.7 Physical intuition of late-time acceleration via VSL
This occurs because the $C$–photons traveled faster at the beginning of their journey toward Earth and must originate from a farther distance (thus appearing fainter than expected) to accumulate sufficient redshift. In either case, the VSL-based Hubble diagram exhibits a steeper upward slope in the high-$z$ section. This intuition is illustrated in Fig. 4.
An alternative explanation for late-time acceleration is a decrease in the speed of light in the intergalactic space as the universe expands. In summary, VSL cosmology can account—qualitatively and quantitatively—for the distance modulus excess observed in high-$z$ SNeIa without the need for dark energy. This approach overcomes the fine-tuning and coincidence problems associated with the concordance $\Lambda$CDM model. Furthermore, it bypasses the enigmatic nature of dark energy and aligns with BDRS’s 2003 groundbreaking analysis of the CMB power spectrum which also negates the dark energy hypothesis [146].
VI.8 A potential resolution to the $H_{0}$ tension via an astronomical origin
In recent years, interest in the $H_{0}$ tension has intensified [1, 2, 150]. One popular approach to address this issue involves treating $H_{0}$ as a “running” value, meaning that its value depends on the redshift of the data used to deduce it; e.g., see Refs. [153, 151, 152].
Our VSL cosmology naturally produces a “running” $H_{0}$ through the function $F(z)$. As mentioned in Section VI.1, $F(z)$ represents the change in the typical size of galaxies as they evolve in an expanding intergalactic space (see also Ref. [19] for clarifications). From the modified formula (58), we can define the “running” value for $H_{0}$ as
@@ VI.8 A potential resolution to the $H_{0}$ tension via an astronomical origin
Using (55) and the value $F_{\infty}=0.93$ obtained from the fit in Section VI.2, the computed “running” $H_{0}(z)$ is depicted as dashed curves in Fig. 3. At the time of recombination, the “running” value for $H_{0}$ would be $H_{0}\,F_{\infty}=47.2\times 0.93=43.9$, representing a $7\%$ reduction from its value estimated using low-redshift SNeIa. The function $F(z)$ hence shed important light on the $H_{0}$ tension. Furthermore, in [147] Hunt and Sarkar obtained $H_{0}\approx 44$ from the CMB power spectrum using a supergravity-based multiple inflation scenario. Our result of $H_{0}(z\gg 1)\approx 43.9$ is thus in agreement with that of the Hunt–Sarkar team.
Importantly, the function $F(z)$ is a well-defined and physically motivated concept: it models the evolution of galaxies in an expanding intergalactic space. Consequently, the “running” $H_{0}$ in our approach is rooted in an astronomical origin, rather than a cosmic origin.
VI.9 An infinite cosmological horizon
In the modified FLRW metric (49), the cosmological horizon at the recombination time $t_{rec}$ is given by
@@ VI.9 An infinite cosmological horizon
VII Conclusions
The action given in Eqs. (7), (8), and (12) with $V(\Phi)\propto\Phi^{4}$ describes a generally covariant theory of gravitation and matter, which also possesses dilatation symmetry. This action is able to evade the observational bounds on the fifth force. A notable feature is the non-minimal coupling of matter with gravity, expressed through the coupling of the Higgs field $\Phi$ with a background dilaton field $\chi$ in the form $\chi^{2}\,\Phi^{2}$. Crucially, the (dimensionful) Planck constant $\hbar$ and speed of light $c$ are absent in this adimensional action. Previously, Fujii and Wetterich utilized this action to generate variable particle masses and Newton’s ‘constant’ [12, 13, 14, 15, 16]. In Sections II and III of this paper, we developed an alternative mechanism to construct (variable) $\hbar$ and $c$ from the background dilaton field $\chi$. Our approach can be generalized when the matter Lagrangian (8) is extended to cover the Glashow–Weinberg–Salam (GWS) model of the $SU(3)\times SU(2)\times U(1)$ gauge group.
In our approach, each open set vicinity of any given point $x^{*}$ on the spacetime manifold is equipped with a replica of the GWS action for matter fields (fermions, gauge vector bosons, and Higgs scalar bosons), operating with its own effective values of $\hbar_{*}$ and $c_{*}$. Lorentz symmetry holds locally within each open set. These GWS replicas are isomorphic (i.e. structurally identical), differing only in the values of $\hbar_{*}$ and $c_{*}$, which are determined by the prevailing value of the background dilaton field $\chi(x^{*})$ per $\hbar_{*}=\chi^{-1/2}(x^{*})$ and $c_{*}=\chi^{1/2}(x^{*})$. Crucially, within each open set, the length and time scales of physical processes are governed by the dilaton field as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$, resulting in an anomalous anisotropic relation between the rate of clocks and the length of rods, $\tau^{-1}\propto l^{-3/2}$. This novel relation enables us to predict a new “time dilation effect of the Third kind”, to be reported elsewhere [154].
Our construction hence fulfills Einstein’s original—albeit long overlooked—vision in incorporating the local validity of Lorentz symmetry, the variation of speed of light in curved spacetimes, and the geometric nature of gravity in a unified framework [21].
The anisotropic relation $\tau^{-1}\propto l^{-3/2}$ mentioned above gives rise to new physics and phenomenology. Section VI of our paper presented key findings on the phenomenological consequences of our mechanism in the context of cosmology, particularly addressing late-time cosmic acceleration via the variation in $c$ while eliminating the need for dark energy. Remarkably, our results revive a crucial—but previously overlooked—discovery made by Blanchard, Douspis, Rowan-Robinson, and Sarkar in 2003 which circumvented the requirement for dark energy in their analysis of the CMB power spectrum [146].
@@ Acknowledgements.
I thank Clifford Burgess, Tiberiu Harko, Robert Mann, Anne-Christine Davis and Eoin Ó Colgáin for their constructive and supportive comments during the development of this work.
————$\infty$—————
@@
In our mechanism, the electron mass $m$ and charge $e$ are dimensionless quantities derived from the parameters $f$, $\mu$, $\lambda$, and $\alpha$ via Eq. (19). The dilaton field $\chi$ is the single entity that carries a scale, which then serves as the only unit available. Consequently, all physical quantities can be expressed exclusively in terms of the magnitude of $\chi$. For instance, $\hbar$ and $c$ have the units of $\chi^{-1/2}$ and $\chi^{1/2}$, respectively, whereas $G$ is dimensionless. The rest mass energy $E=mc^{2}$ has the unit of $\chi$, owing to $m$ being dimensionless and $c\propto\chi^{1/2}$. Similarly, the Coulomb potential of a point charge $e$, given by $e^{2}/r$, also carries the unit of $\chi$, by virtue of $e$ being dimensionless and $r\propto\chi^{-1}$. In general, a theory with dilatation symmetry would require only one ‘unit’ (associated with the dilaton field $\chi$), rendering the canonical three-unit system (mass $M$, length $L$, time $T$) unnecessary.
@@
Our new unit system based on the trio $\{e_{0},\,m_{0},\,\chi_{*}\}$ deviates from the Planck $cGh$ unit system, which relies on the trio of constants $\{c,\,G,\,\hbar\}$. Note that our new unit system is akin to the unit system introduced by Hartree in 1928 [155], based on the electron charge $e$, mass $m$, and the Bohr radius $a_{B}$ (the typical size of the hydrogen atom in its ground state). In the Hartree system, the Planck constant and speed of light are defined via $\{e,\,m,\,a_{B}\}$ as
$\hbar$ $=e\sqrt{m\,a_{B}};\ \ \ c=\frac{e}{\alpha\sqrt{ma_{B}}}$ (76)
in close resemblance to the first two identities in Eq. (75).